OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)?

sam (.sam.):

first find the slope, you know the slope formula?

OpenStudy (anonymous):

if you could instruct just a little id appriciate it...

sam (.sam.):

\[slope=\frac{y-y_1}{x-x_1}\] \[slope=\frac{-1-(-7)}{5-2}\]

OpenStudy (anonymous):

and would that be the final answer?

sam (.sam.):

so, you'll get 2 for slope, then use y-y1=m(x-x1) pick any points (2,-7) y-(-7)=2(x-2) y+7=2x-4 add both sides by -7 and you're done

OpenStudy (anonymous):

so add 7 to the one above?

sam (.sam.):

no , add -7

OpenStudy (anonymous):

oh ok thanks so much!

sam (.sam.):

welcome!